OpenStudy (anonymous):

Evaluate the expression when m = 2. 11 3m (Points : 1) 5 6 10 12

OpenStudy (anonymous):

@AllTheLonelyKillers

OpenStudy (zzr0ck3r):

\[11\times 3M\]?

OpenStudy (anonymous):

whats the sin in between ?

OpenStudy (zzr0ck3r):

must be -

OpenStudy (zzr0ck3r):

\[11-3m\]?

OpenStudy (anonymous):

oh sorry - sign

OpenStudy (zzr0ck3r):

what is 3*2?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

whats in between the 11 and the 3

OpenStudy (akashdeepdeb):

All you need to do is SUBSTITUTE the m value which is given to you as m = 3 in the expression 11 - 3m = 11 - 3*2 = 11 - 6 Understood this? :)

OpenStudy (zzr0ck3r):

what is 11-6?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

5

OpenStudy (zzr0ck3r):

correct

OpenStudy (akashdeepdeb):

And that is your answer! :D

OpenStudy (anonymous):

lol u guys good idk who to give the medal 2

OpenStudy (akashdeepdeb):

zzrocker.